# Attack Anything: Blind DNNs via Universal Background Adversarial Attack

### v1

@@ IV-E3 Across Different Tasks
IV-F Ablation Study
To verify the effectiveness of the proposed ensemble attack strategy, we compare the attack performance of the ensemble attack with the single attack. The experimental results are shown in Table. VI. It is observed that the attack transferability is significantly improved by the proposed ensemble strategy.
To verify the key value of smoothness loss for physical attacks, we compare the attack efficacy of smooth and smooth-less perturbations under various thresholds of confidence score as shown in Table VII. The experimental results demonstrate that smoothness loss plays an indispensable role in conducting physical attacks.
V Discussion

### v3

@@ 4.5.3 Across Different Tasks
4.6 Ablation Study
To verify the effectiveness of the proposed ensemble attack strategy, we compare the attack performance of the ensemble attack with the single attack. The experimental results are shown in Table. 5. It is observed that the attack transferability is significantly improved by the proposed ensemble strategy.
To verify the key value of smoothness loss for physical attacks, we compare the attack efficacy of smooth and smooth-less perturbations under various thresholds of confidence score as shown in Table 7. The experimental results demonstrate that smoothness loss plays an indispensable role in conducting physical attacks.
5 Discussion